OpenStudy (kl0723):

Integral:

OpenStudy (kl0723):

\[\int\limits \arctan4tdt\]

OpenStudy (amistre64):

any ideas?

OpenStudy (kl0723):

maybe set u=4t and dv=arctan?

OpenStudy (amistre64):

arctan is the name of a function, so i dont think you can do that ....

OpenStudy (kl0723):

:/ I need to approach by integration by parts... but I'm a little confused with this one

OpenStudy (amistre64):

let u=t

OpenStudy (amistre64):

might have the uv parts backwards u = arctan(4t) dv = t might be better

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=integrate+arctan%284t%29+dt this would suggest a by parts setup

OpenStudy (amistre64):

since arctan 4t is in the left term, and we are trying to integrate it ... id say its best to let that be u v = t, so dv = 1

OpenStudy (amistre64):

u = arctan(4t) v = t du = dv = 1 tan u = 4t sec^2u du = 4 dt du = 4 dt/(tan^2u + 1) = 4dt/(16t^2+1)

OpenStudy (amistre64):

so that simplifies it quit alot:\[\int tan^{-1}(4t)dt=t~tan^{-1}(4t)-\int\frac{4t}{16t^2+1}dt\]

OpenStudy (kl0723):

how do you get 16t^2 + 1 right there?